An auditor of Health Maintenance Services of Georgia reports 40 percent of policyholders 55 years or older submit a claim during the year. Fifteen policyholders are randomly selected for company records.
A. How many of the policy holders would you expect to have filed a claim within the last year?
B. What is the probability that 10 of the selected policyholders submitted a claim last year?
C. What is the probability that 10 or more of the selected policyholders submitted a claim last year?
D. What is the probability that more than 10 of the selected policyholders submitted a claim last year?

Answer by stanbon(75887)

An auditor of Health Maintenance Services of Georgia reports 40 percent of policyholders 55 years or older submit a claim during the year. Fifteen policyholders are randomly selected for company records.
A. How many of the policy holders would you expect to have filed a claim within the last year?
----
expected value = np = 15*0.4 = 6
-----------------------------------------------
B. What is the probability that 10 of the selected policyholders submitted a claim last year?
P(x = 10) = 15C10*0.4^10*0.6^5 = binompdf(15,0.4,10) = 0.0245
-----------------------------------------------------------
C. What is the probability that 10 or more of the selected policyholders submitted a claim last year?
P(10<= x <= 15) = 1 - binomcdf(15,0.4,9) = 0.0338
-----------------------------------------------------------
D. What is the probability that more than 10 of the selected policyholders submitted a claim last year?
P(11<= x <=15) = 1 - binomcdf(15,0.4,10) = 0.0093
==============================
Cheers,
Stan H.
==============================